OpenStudy (ckenn):

A cube has a volume of 36 cubic inches. What is the volume of a pyramid that fits exactly inside the cube?

OpenStudy (ckenn):

@Directrix could you help me? (:

OpenStudy (ckenn):

@johnweldon1993 @welshfella (:

Directrix (directrix):

Are you given that the pyramid is regular?

OpenStudy (ckenn):

no

OpenStudy (ckenn):

@Ms-Brains

Directrix (directrix):

Courtesy of @campbell_st Volume of a Pyramid is V = 1/3 length x width x height Volume of a cube is V = length x width x height

Directrix (directrix):

volume of pyramid = 1/3 of cube =(1/3)36 =12 cubic feet

Directrix (directrix):

Read the entire thread here: http://openstudy.com/updates/542b1dc1e4b0173657d9c200

OpenStudy (ckenn):

thank you

Directrix (directrix):

You are welcome.

MsBrains (ms-brains):

Hey Buddy, you still need my help?